OpenStudy (anonymous):

Write Each Number Without An Exponent: 6.021967*10^5 p.s I need to show all work

OpenStudy (dan815):

hello I would like to help you!!

OpenStudy (dan815):

my friend these numbers are a beautiful thing you see

OpenStudy (dan815):

you see we write our numbers in a decimal system deci = 10 that means that for every 10 units we add another digit like 0-9 are 1 digit but when its 10 it means ---,---- 1 , 0 we have 1 in the 2nd place holder and 0 in the first place holder when we have the number 12 that means we have ----,---- 1 , 2 1 in the 2nd place holder and 2 in the first place holder so we have 1 of those 10s and then 2 of those 1s each holder is defined liked this 10^0 = your 1s 10^1 = your 10s 10^2 = your 100s similarily 10-1 = 1/10 10-2 = 1/100 the decimal is like a place holder if you multiply from 10 then you are moving the decimal place to the right by 1 if u multiply by 10 twice so 10*10 = 10*2 you will move the decimal place holder 2 times to the right

OpenStudy (dan815):

say i have a number 10.0002 * 10^1 or 10 we will move place holder once to the right 100.002 10.0002* 10^2 = 1000.02 we have moved the decimal 2 times 10.0002* 10^3=10000.2 we have moved it 3 times 10.0002* 10^4 = 100002. we have moved it 4 times there are endless number of 0s behind the decimal but we dont write it for convenience 100002.00000 we can go on and on and on